Electrostatics

a particle of mass m and charge q is placed in an electric field e which varies with time t as e=e0sinwt. it will undergo shm of amplitude?

sagar singh
13 Years agoGrade

2 Answers

Prajwal kr
13 Years ago
F=qe0sinwt. a= (qesinwt)/m dv=((qesinwt)/m) dt Integrating: V= - (qecoswt)/wm dx/dt= - (qecoswt)/wm Integrating the same way: X= - (qesinwt)/w2m Comparing with x=Asinwt; A=(qe0)/w2m We can ignore the minus as it only signifies initial direction and nothing more than that.
Kushagra Madhukar
6 Years ago
Dear student,
Please find the solution to your problem.
 
F = qe0sinwt.
a = (qesinwt)/m
dv = ((qesinwt)/m) dt
 
Integrating:
v = – (qecoswt)/wm
dx/dt = – (qecoswt)/wm
 
Integrating the same way:
x = – (qesinwt)/w2m
 
Comparing with x = Asinwt;
A = (qe0)/w2m
(We can ignore the minus as it only signifies initial direction and nothing more than that)
 
Thanks and regards,
Kushagra